OpenStudy (anonymous):

Convert the following equation from point slope to slope intercept form. In other words, if the equation is rewritten to look like y=mx+b, what are the values of m and b? y+15=4(x+3)

OpenStudy (anonymous):

Well you want y+15=4(x+3) to look like y=mx+b you want to isolate y so how do you do that?

sam (.sam.):

y+15=4(x+3) y+15=4x+12 y=-3+4x m=4 y-intercept=-3

OpenStudy (vishal_kothari):

y=4x-3

OpenStudy (anonymous):

w/e just give them the answers

Directrix (directrix):

y+15=4(x+3) y = 4(x+3) - 15 y = 4x + 12 - 15 y = 4x -3 y = 4x + (-3) slope = 4 or 4/1 y-intercept = -3